OpenStudy (anonymous):

Prove that in any triangle the length of a side is less than half of the perimeter.

OpenStudy (anonymous):

Lets consider a Triangle with area greater than Zero (obviously :/) \[\sqrt{s(s-a)(s-b)(s-c)} = Area \implies Area^2 = s(s-a)(s-b)(s-c) \] If Area > 0 \(\implies \) \(Area^2 > 0\). \[s(s - a)(s-b)(s-c)> 0 \] Note. s is the half perimeter.

OpenStudy (anonymous):

I have never ever done this problem before ...So ... DO NOT TRUST ME.

OpenStudy (anonymous):

Lets not square it, okay? \[\sqrt{s(s-a)(s-b)(s-c)} > 0\] s = (a+b+c)/2 \[\sqrt{2s(2s-2a)(2s-2b)(2s-2c)} > 0\]\[\sqrt{(a+b+c)(b + c - a)(a+b-c)(a+c - b)} >0\] For a triangle, third side can not be greater than the sum of the other two sides.

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

yes it is

OpenStudy (amistre64):

what does area have to di with perimeter?

OpenStudy (amistre64):

id prove it for an equitri and asses it from there

OpenStudy (amistre64):

i see you used heron for relating sides ...

OpenStudy (anonymous):

a+b>c a+b+c>2c (a+b+c)/2 > c

OpenStudy (anonymous):

pssstt... i am still not sure if it's a conclusive proof, just made it out of thin air :/

OpenStudy (anonymous):

oh foolformath completed it

OpenStudy (amistre64):

how many assumptions are we allowed to make? fool used a thrm for a proposition i think

OpenStudy (anonymous):

a + b > c a+ b - c + c -c>0 2s - 2c > 0 s > c same can be done for every side

OpenStudy (anonymous):

Sum of any two sides of a triangle is greater than the third side. http://www.proofwiki.org/wiki/Sum_of_Two_Sides_of_Triangle_Greater_than_Third_Side